OpenStudy (anonymous):

Find the values of a and b that make f continuous everywhere: f(x) =(x^2 -4)/x-2, x<2 f(x) =ax^2 +bx +3, 2= 3

myininaya (myininaya):

this looks like a problem for my paper lol

OpenStudy (anonymous):

what school do you go to?

myininaya (myininaya):

scanning i don't go anymore

myininaya (myininaya):

see we have two equations 8a+2b=-1 10a+2b=3

myininaya (myininaya):

if we solve the first one for b, we get 2b=-1-8a b=(-1-8a)/2 then we can plug this into the other equation where the other b is 10a+2(-1-8a)/2=3 10a+(-1-8a)=3 10a-1-8a=3 2a-1=3 2a=4 a=2 so b=(-1-8*2)/2=(-1-16)/2=-17/2

myininaya (myininaya):

cool? :)

OpenStudy (anonymous):

yeah i think so. Thanks a lot

myininaya (myininaya):

np i like the continuous problems they are fun

OpenStudy (anonymous):

I think you made a slight mistake in the first equation. It should be \(4a+2b+3=4\). Right?!

OpenStudy (anonymous):

So, we have the two equations \(4a+2b=1\) and \(10a+2b=3\). Subtract equation (1) from equation (2), you get \(6a=2 \implies a=\frac{1}{3}\). Substitute in either equation, you will find that \(b=-\frac{1}{6}\).